OpenStudy (anonymous):

(sinx/1-cosx)-(sinxcosx/1+cosx)=cscx(1+cos^2x)

OpenStudy (anonymous):

RHS = 1+ cos x LHS = (sinx ( 1+ cos x ) - (sinx*cos^2x))/sin^2 x simplify you get (cos x - sin^3x)/sin^2x cos x - sin^3 x = sin^2 x What are we to do exactly? And i did that mentally so it might be wrong just check it please :P

OpenStudy (anonymous):

Sorry we are supposed to verify that each trigonometric equation is an identity.

OpenStudy (anonymous):

Oh... doing this mentally is so much harder than i thought =.= See i have a simple trick... What you do is simplify every trigonometric notion into simplest form, and keep using identities. here the only one which is of any use is sin^2 + cos^2 = 1 So you keep using that and if you cant ask me ill try again :)

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

Numerator of the left side of the identity: sin x(1 + cos x)- sin x*cos x(1 - cos x) = = sin x + sin x*cos x - sin x*cos x + sin x cos^2 x = sin x(1+ cos ^2x) Denominator: (1 - cos x)(1 + cos x) = 1 - cos^2 x = sin^2 x Left side after simplification: (1 + cos^2 x)/sin x = csc x(1 + cos ^2 x) That proves the identity.

OpenStudy (anonymous):

^ Perfect :)